OpenStudy (anonymous):

Dividing Polynomial question Find the missing dividend or divisor ?/x-5=x-11

myininaya (myininaya):

How would you solve this for y? y/2=8

OpenStudy (anonymous):

Try and do it and I will learn from the steps ,I have done some other problem in polynomial but i dont understand when they ask find the missing dividend or divisor thats why i post it here

myininaya (myininaya):

\[\frac{y}{2}=8\] Since we want y and y is being divided by 2, to undo the division by 2 we multiply by 2. If we multiply by 2 on one side then we do it to the other side as well. \[2 \cdot \frac{y}{2}=2 \cdot 8\] \[y=16\]

myininaya (myininaya):

Now and your question you are asked to find the ? in: \[\frac{?}{x-5}=x-11 \] You do the same thing we did to find y.

myininaya (myininaya):

If we have p/q, then q is the divisor and p is the dividend.

OpenStudy (anonymous):

so is 6

myininaya (myininaya):

No...

myininaya (myininaya):

Like when we had y/2=8 To find y we just multiply both sides by what it is being divided by. So do that here for the ?.

OpenStudy (anonymous):

I will figure it out myself ,I appreciate thanks

myininaya (myininaya):

If you have ?/(x-2)=x-1 All you do is multiply both sides by (x-2) like this ?=(x-2)(x-1)

myininaya (myininaya):

Just like if you had y/8=2 multiply both sides by 8 y=8(2)

myininaya (myininaya):

If you had ?/star=fish, then ?=star*fish

OpenStudy (anonymous):

Dont worry anymore .I will figure the problem myself

myininaya (myininaya):

I'm only worried you are making this way harder than it is.

OpenStudy (anonymous):

I am not ,I ask you if you can help me solved it ,Math is hard ,I was given 30 questions on polynomial and I have already solved 28 questions ...I only have Just 2 question that I think is given me headache ..So I will relax a little and I will go back to it and do it myself ...I was thinking I can see someone to help me with it .

myininaya (myininaya):

So like I'm trying to be helpful here I promise. Like tell me what you don't understand on how to get the ? by itself? \[\frac{?}{x-5}=x-11\]

myininaya (myininaya):

Or what you don't understand about the explanation I have given?

myininaya (myininaya):

To undo division by (x-5), we multiply both sides by what